Differentiate (x5 + 6)5 using chain rule
1
Expert's answer
2020-09-14T14:51:37-0400

"f(x)=(x^5+6)^5" .

"f'(x)=5(x^5+6)^4*\frac{d}{dx}(x^5+6)=5(x^5+6)^4*5x^4=25x^4(x^5+6)^4."